# InterMask: 3D Human Interaction Generation via Collaborative Masked Modelling

### v1

@@ Abstract
Generating realistic 3D human-human interactions from textual descriptions remains a challenging task. Existing approaches, typically based on diffusion models, often generate unnatural and unrealistic results. In this work, we introduce InterMask, a novel framework for generating human interactions using collaborative masked modeling in discrete space. InterMask first employs a VQ-VAE to transform each motion sequence into a 2D discrete motion token map. Unlike traditional 1D VQ token maps, it better preserves fine-grained spatio-temporal details and promotes spatial awareness within each token. Building on this representation, InterMask utilizes a generative masked modeling framework to collaboratively model the tokens of two interacting individuals. This is achieved by employing a transformer architecture specifically designed to capture complex spatio-temporal interdependencies. During training, it randomly masks the motion tokens of both individuals and learns to predict them. In inference, starting from fully masked sequences, it progressively fills in the tokens for both individuals. With its enhanced motion representation, dedicated architecture, and effective learning strategy, InterMask achieves state-of-the-art results, producing high-fidelity and diverse human interactions. It outperforms previous methods, achieving an FID of $5.154$ (vs $5.535$ for in2IN) on the InterHuman dataset and $0.399$ (vs $5.207$ for InterGen) on the InterX dataset. Additionally, InterMask seamlessly supports reaction generation without the need for model redesign or fine-tuning.
@@ 1 Introduction
Existing works on human interaction generation commonly rely on diffusion models. ComMDM (Shafir et al., 2024) bridges two pretrained single-person motion diffusion models using a small neural layer with limited interaction data. In contrast, InterGen (Liang et al., 2024) proposes a specifically tailored diffusion model for two-person interactions, which uses cooperative transformers to model each individual’s motion conditioned on the latent features of their partner’s motion. in2IN (Ruiz-Ponce et al., 2024) then extends it by conditioning the diffusion model on LLM-generated individual descriptions in addition to the overall interaction descriptions, and leveraging motion priors learned from single-person motion datasets like HumanML3D (Guo et al., 2022a). MoMat-MoGen Cai et al. (2024) retrieves motions from supplementary datasets, based on textual inputs, and refines them for interaction quality. Some very recent efforts (Shan et al., 2024; Fan et al., 2024) extend these methods to more than two individuals. Despite these efforts, the generated two-person interactions still fall short from achieving satisfactory realism and fidelity.
@@
In the second stage, a specialized Inter-M Transformer (section 3.2) is trained to collaboratively model the tokens of both individuals using a generative masked modeling framework. During training, a random proportion of tokens from either both individuals or only one individual is masked and predicted. The masking proportion follows a scheduling function to ensure its tractability during inference. The inference process begins with all tokens masked, which are progressively generated over a pre-defined number of iterations. At each iteration, the model predicts all masked tokens, but only the most confident predictions are retained, while the rest are re-masked and re-predicted in subsequent iterations. In addition to the standard self attention module, Inter-M Transformer features a shared spatio-temporal attention module to capture fine-grained dependencies within each individual’s motion, and a cross-attention module to capture dependencies between the two individuals’ motion. This design effectively the dynamics of spatial and temporal relationships with and between the two individuals, allowing enhanced interaction generation capabilities.
The main contributions of this work are as follows. First, we introduce InterMask, a novel masked generative framework for human-human interaction generation from textual descriptions. InterMask utilizes a 2D VQ encoding that transforms motion sequences into discrete 2D token maps and features a dedicated transformer trained with a generative masked modeling framework which effectively captures both intra- and inter-person spatial-temporal dependencies. Second, InterMask achieves state-of-the-art results in generating high-fidelity, text-conditioned human interactions. Empirically, it achieves an FID of 5.15 (vs. 5.54 in Ruiz-Ponce et al. (2024)) on InterHuman and 0.399 (vs. 5.207 in Liang et al. (2024)) on InterX. Third, InterMask seamlessly supports tasks such as reaction generation without requiring any task-specific fine-tuning or architectural changes.
@@ 3 Methodology
We create a 2D token map with both spatial and temporal dimensions by representing motion as $\mathbf{m}_{p}\in\mathbb{R}^{N\times J\times d}$ and using 2D convolutions. As illustrated in Figure 2(a), our VQ-VAE downsamples the input sequence $\mathbf{m}_{p}$ into a latent representation $\mathbf{\tilde{t}_{p}}\in\mathbb{R}^{n\times j\times d^{\prime}}$, with downsampling ratios $n/N$ and $j/J$. Each $d^{\prime}$-dimensional vector is then quantized by replacing it with the nearest entry from a learnable codebook $\mathcal{C}=\{\mathbf{c}_{k}\}_{k=0}^{|\mathcal{C}|-1}\subset\mathbb{R}^{d^{\prime}}$, producing a quantized sequence $\mathbf{t}_{p}=\mathrm{Q}(\mathbf{\tilde{t}}_{p})\in\mathbb{R}^{n\times j\times d^{\prime}}$. This is then decoded to reconstruct the motion $\mathbf{\hat{m}_{p}=\mathrm{D}(\mathbf{t}_{p})}$. After training, each vector in $\mathbf{t}_{p}$ can be replaced by its respective index $k$ in $\mathcal{C}$ to obtain the 2d discrete representation of motion, namely motion tokens $\{t_{p}\}_{p\in\{a,b\}}\in\{0,1,\cdots,|\mathcal{C}|-1\}^{n\times j}$, which is modeled in our Inter-M Transformer.
@@ 3.1.1 Training Objective
We also incorporate geometric losses from Liang et al. (2024) to impose additional constraints, as shown in eq. 2. The velocity loss $\mathcal{L}_{vel}$ encourages the reconstructed motion sequences to obey the velocity of joints in the ground truth sequences and the bone length loss $\mathcal{L}_{bl}$ the distance between adjacent joints. The foot contact loss $\mathcal{L}_{f}c$ encourages the feet to have zero velocity whenever they are in contact with the ground.
The overall loss function $\mathcal{L}_{vqvae}$ is a weighted sum of the primary VQ loss and geometric losses:
@@ 3.2.1 Token Masking and Embedding
In the second stage, we employ step unroll masking, adopted from Kim et al. (2023b), where we remask a portion ($\gamma(\tau_{i+1})$) of the predicted tokens with the lowest confidence scores and predict them again. This refines the predictions iteratively, aligning with the inference process, where tokens are progressively revealed. The masked tokens $\tilde{t}$ are then embedded using the input process module, consisting of a learnable embedding layer and a linear transformation. Then, 2D positional encodings, from  Wang & Liu (2021), are added to impose spatio-temporal structure on the embeddings. The updated embeddings, $\mathbf{e}^{0}\in\mathbb{R}^{(2nj+1)\times\tilde{d}}$, where $\tilde{d}$ is the transformer embedding dimension, are then passed through $\mathbf{L}$ blocks of the Inter-M transformer.
@@ 4 Experiments
Datasets We adopt two datasets to evaluate InterMask for the text-conditioned human interaction generation task: InterHuman (Liang et al., 2024) and InterX (Xu et al., 2024a). The InterHuman dataset contains 7,779 interaction sequences and InterX contains 11,388, each paired with 3 distinct textual annotations. InterHuman follows the AMASS (Mahmood et al., 2019) skeleton representation with 22 joints, including the root joint. Each joint is represented by $\{p_{g},v_{g},r_{6d}\}$, where $p_{g}\in\mathbb{R}^{3}$ is the global position, $v_{g}\in\mathbb{R}^{3}$ is the global velocity, and $r_{6d}\in\mathbb{R}^{6}$ is the local 6D rotation of each joint, rendering $\mathbf{m}_{p}\in\mathbb{R}^{N\times 22\times 12}$. InterX follows the SMPL-X (Pavlakos et al., 2019) skeleton representation, comprising 54 body, hands and face joints, accompanied by root orientation and translation. Each joint and root orientation is represented by $\{r_{6d}\}$ and root translation by $\{p_{g}\}$, to which we incorporate root velocity rendering root $\{p_{g},v_{g}\}$ and $\mathbf{m}_{p}\in\mathbb{R}^{N\times 56\times 6}$. We adhere to the respective body representations of both datasets to demonstrate that our framework is compatible with different representation formats and joint counts. Other implementation details including architecture, training and inference hyper-parameters are provided in appendix A.
@@ 4.1 Comparison with state-of-the-art approaches
Quantitative Comparison Table 1 shows quantitative comparison of our InterMask with previous methods. All evaluations are run 20 times (with the exception of MModality, which is run 5 times), and we report the averaged results along with a 95% confidence interval. InterMask achieves state-of-the-art (SOTA) results on both InterHuman and InterX datasets. It records the lowest FID scores (5.154 on InterHuman and 0.399 on InterX), indicating superior realism and quality of generated interactions, and leads in R-Precision and MMdist, showing excellent text alignment. With different body representations in each dataset, these results highlight the robustness of our method, proving it is not dependent on a specific body pose structure. Furthermore, the larger performance gap on the larger InterX dataset suggests that InterMask scales effectively with more data. While our MModality is slightly lower than some methods, the high R-Precision and low MMdist emphasize that InterMask prioritizes adherence to text over extreme diversity, yet still achieves high Diversity scores, demonstrating its ability to generate a broad range of interactions. Please refer to appendix B for visual demonstration of InterMask’s ability to generate diverse results from the same text prompt.
Qualitative Comparison In Figure 4, we provide a qualitative comparison of interaction sequences generated by our InterMask and InterGen Liang et al. (2024), trained on the InterHuman dataset, for the same text descriptions. For the prompt ”The first person raises the right leg aggressively towards the second”, InterGen incorrectly raises both legs, while InterMask accurately raises only the right leg. For ”Two people bow to each other”, InterGen introduces an unnecessary fighting gesture after the bow, indicating overfitting to implicit biases in the training data, whereas InterMask stays true to the simple bow. Lastly, for ”One person is sitting and waving their hands at the other person, while the other drifts away”, InterGen generates a crouching pose instead of sitting and misplaces the waving action, while InterMask faithfully generates the described scene. These examples demonstrate that InterMask produces more higher quality and more reliable interactions compared to InterGen.
@@
User Study User studies are essential in generative AI, complementing metrics like FID by incorporating human perception to evaluate the realism and quality of generated content. To compare our InterMask with InterGen (Liang et al., 2024), we conducted a user study where 30 interaction sequences were generated by both models with identical motion lengths. Each sequence pair was rated by 3 users on interaction quality and adherence to text (scores out of 5), and users selected their preferred sequence based on both factors. More details are provided in appendix F. As shown in Figure 5, InterMask outperforms InterGen in interaction quality (4.089 vs. 3.296) and text adherence (3.938 vs. 3.198), with 69.14% of users preferring InterMask over InterGen.
@@ 4.2 Ablation Studies
Figure 7: Qualitative results for the ablation study on Inter-M Transformer to verify the proposed Spatio-Temporal Attention module.
Motion VQ-VAE Table 2 presents an ablation study to evaluate key components of our proposed Motion VQ-VAE. We investigated two primary factors: the input modality (individual motions vs two-person interactions) and the token map dimensions (1d vs 2d), using Reconstruction FID and MPJPE (Mean Per Joint Position Error). The results demonstrate that using individual motions as input with a 2D token map outperforms both the baseline of individual motions with a 1D token map, and the two-person interaction with a 2D token map. These findings suggest that preserving spatial information through 2D token maps and focusing on individual motions performs accurate motion reconstruction, and creates a more expressive token map with high-fidelity movement details. This can be further verified by the qualitative comparison provided in Figure 6. As shown, the 1D VQ-VAE struggles to accurately reconstruct the spatial positions and orientations of the joints for both individuals, leading to not only unrealistic interactions but also abnormal individual poses. In contrast, the proposed 2D VQ-VAE provides highly accurate reconstruction at both the individual pose and the collective interaction level.
Inter-M Transformer Table 3 presents an ablation study to evaluate key components of our Inter-M Transformer. We compared collaborative modeling against alternative modeling and examined the impact of various attention mechanisms and masking strategies. Alternative Modeling (appendix C) models tokens of one person at a time, while conditioning in on the tokens of the other person, leading to an alternative generation during inference. Collaborative Modeling on the other hand, models tokens of both individuals simultaneously. Our full InterMask model with collaborataive modeling outperforms all ablated versions, particularly in R-Precision and FID scores. Notably, removing any of the attention mechanisms or employing alternative modeling severely degraded performance, with spatio-temporal attention being the most crucial one. Additionally, step unroll and interaction masking proved to be further fine-tuning the results. These results highlight the synergistic importance of all components in our transformer architecture for effective interaction modeling. The contribution of spatio-temporal attentionis further verified by the qualitative comparison illustrated in Figure 7. While the model without spatio-temporal attention generates somewhat plausible individual poses, such as punching and blocking in the boxing scenario, it fails to fully capture the dynamics of the interaction, leading to incorrect positioning between the two individuals. Similarly, in the sneaking scenario, the model without spatio-temporal attention does not capture the gradual change in distance between the two individuals, where one should be moving closer to the other. In contrast, the model with spatio-temporal attention accurately captures these spatial dynamics and interactions.
@@ 4.3 Application: Reaction Generation
One of the key strengths of our InterMask model is its seamless support for the task of reaction generation with or without a text condition, where the motion of one person in an interaction is generated based on the provided reference motion of the other person. Unlike previous diffusion-based models, such as InterGen (Liang et al., 2024), InterMask can perform this task without any task-specific fine-tuning or architectural re-design. This flexibility arises from our use of the masked modeling technique combined with an individual motion tokenizer, which does not require interactions as input. During inference, we simply tokenize the reference motion, leave its tokens unmasked, and progressively generate tokens of the other person. More details are provided in appendix D.

### v3

# InterMask: 3D Human Interaction Generation via Collaborative Masked Modeling

@@ Abstract
Generating realistic 3D human-human interactions from textual descriptions remains a challenging task. Existing approaches, typically based on diffusion models, often produce results lacking realism and fidelity. In this work, we introduce InterMask, a novel framework for generating human interactions using collaborative masked modeling in discrete space. InterMask first employs a VQ-VAE to transform each motion sequence into a 2D discrete motion token map. Unlike traditional 1D VQ token maps, it better preserves fine-grained spatio-temporal details and promotes spatial awareness within each token. Building on this representation, InterMask utilizes a generative masked modeling framework to collaboratively model the tokens of two interacting individuals. This is achieved by employing a transformer architecture specifically designed to capture complex spatio-temporal inter-dependencies. During training, it randomly masks the motion tokens of both individuals and learns to predict them. For inference, starting from fully masked sequences, it progressively fills in the tokens for both individuals. With its enhanced motion representation, dedicated architecture, and effective learning strategy, InterMask achieves state-of-the-art results, producing high-fidelity and diverse human interactions. It outperforms previous methods, achieving an FID of $5.154$ (vs $5.535$ of in2IN) on the InterHuman dataset and $0.399$ (vs $5.207$ of InterGen) on the InterX dataset. Additionally, InterMask seamlessly supports reaction generation without the need for model redesign or fine-tuning.
[…]
gohar-malik.github.io/intermask
@@ 1 Introduction
Existing works on human interaction generation commonly rely on diffusion models. ComMDM (Shafir et al., 2024) bridges two pretrained single-person motion diffusion models using a small neural layer with limited interaction data. In contrast, InterGen (Liang et al., 2024) proposes a specifically tailored diffusion model for two-person interactions, which uses cooperative transformers to model each individual’s motion conditioned on the latent features of their partner’s motion. in2IN (Ruiz-Ponce et al., 2024) then extends it by conditioning the diffusion model on LLM-generated individual descriptions in addition to the overall interaction descriptions, and leveraging motion priors learned from single-person motion datasets like HumanML3D (Guo et al., 2022a). MoMat-MoGen Cai et al. (2024) retrieves motions from supplementary datasets, based on textual inputs, and refines them for interaction quality. Some very recent efforts (Shan et al., 2024; Fan et al., 2024) extend these methods to more than two individuals. Despite these efforts, the generated two-person interactions still fall short of achieving satisfactory realism and fidelity.
@@
In the second stage, a specialized Inter-M Transformer (section 3.2) is trained to collaboratively model the tokens of both individuals using a generative masked modeling framework. During training, a random proportion of tokens from either both individuals or only one individual is masked and predicted. The masking proportion follows a scheduling function to ensure its tractability during inference. The inference process begins with all tokens masked, which are progressively generated over a pre-defined number of iterations. At each iteration, the model predicts all masked tokens, but only the most confident predictions are retained, while the rest are re-masked and re-predicted in subsequent iterations. In addition to the standard self attention module, Inter-M Transformer features a shared spatio-temporal attention module to capture fine-grained dependencies within each individual’s motion, and a cross-attention module to capture dependencies between the two individuals’ motion. This design effectively models the dynamics of spatial and temporal relationships among and between the two individuals, allowing enhanced interaction generation capabilities.
The main contributions of this work are as follows. First, we introduce InterMask, a novel masked generative framework for human-human interaction generation from textual descriptions. InterMask utilizes a 2D VQ encoding that transforms motion sequences into discrete 2D token maps and features a dedicated transformer trained with a generative masked modeling framework which effectively captures both intra- and inter-person spatial and temporal dependencies. Second, InterMask achieves state-of-the-art results in generating high-fidelity, text-conditioned human interactions. Empirically, it achieves an FID of 5.15 (vs. 5.54 of Ruiz-Ponce et al. (2024)) on InterHuman and 0.399 (vs. 5.207 of Liang et al. (2024)) on InterX. Third, InterMask seamlessly supports tasks such as reaction generation without requiring any task-specific fine-tuning or architectural changes.
@@ 3 Methodology
We create a 2D token map with both spatial and temporal dimensions by representing motion as $\mathbf{m}_{p}\in\mathbb{R}^{N\times J\times d}$ and using 2D convolutions. As illustrated in Figure 2(a), our VQ-VAE downsamples the input sequence $\mathbf{m}_{p}$ into a latent representation $\mathbf{\tilde{t}_{p}}\in\mathbb{R}^{n\times j\times d^{\prime}}$, with downsampling ratios $n/N$ and $j/J$. Each $d^{\prime}$-dimensional vector is then quantized by replacing it with the nearest entry from a learnable codebook $\mathcal{C}=\{\mathbf{c}_{k}\}_{k=0}^{|\mathcal{C}|-1}\subset\mathbb{R}^{d^{\prime}}$, producing a quantized sequence $\mathbf{t}_{p}=\mathrm{Q}(\mathbf{\tilde{t}}_{p})\in\mathbb{R}^{n\times j\times d^{\prime}}$. This is then decoded to reconstruct the motion $\mathbf{\hat{m}_{p}=\mathrm{D}(\mathbf{t}_{p})}$. After training, each vector in $\mathbf{t}_{p}$ can be replaced by its respective index $k$ in $\mathcal{C}$ to obtain the 2d discrete representation of motion, namely motion tokens $\{t_{p}\}_{p\in\{a,b\}}\in\{0,1,\cdots,|\mathcal{C}|-1\}^{n\times j}$, which is modeled in our Inter-M Transformer. More details are provided in appendix B.
@@ 3.1.1 Training Objective
We also incorporate geometric losses from Liang et al. (2024) to impose additional constraints, including the joint velocity loss $\mathcal{L}_{vel}$, the foot contact loss $\mathcal{L}_{fc}$, and the bone length loss $\mathcal{L}_{bl}$, which are explained in appendix C. The overall loss function $\mathcal{L}_{vqvae}$ is a weighted sum of these losses:
@@ 3.2.1 Token Masking and Embedding
In the second stage, we employ step unroll masking, adopted from Kim et al. (2023b), where we remask a portion ($\gamma(\tau_{i+1})$) of the predicted tokens with the lowest confidence scores and predict them again. This refines the predictions iteratively, aligning with the inference process, where tokens are progressively revealed. More details of the masking strategy are provided in appendix D. The masked tokens $\tilde{t}$ are then embedded using the input process module, consisting of a learnable embedding layer and a linear transformation. Then, 2D positional encodings, from  Wang & Liu (2021), are added to impose spatio-temporal structure on the embeddings. The updated embeddings, $\mathbf{e}^{0}\in\mathbb{R}^{(2nj+1)\times\tilde{d}}$, where $\tilde{d}$ is the transformer embedding dimension, are then passed through $\mathbf{L}$ blocks of the Inter-M transformer.
@@ 4 Experiments
Datasets We adopt two datasets to evaluate InterMask for the text-conditioned human interaction generation task: InterHuman (Liang et al., 2024) and InterX (Xu et al., 2024a). The InterHuman dataset contains 7,779 interaction sequences and InterX contains 11,388, each paired with 3 distinct textual annotations. InterHuman follows the AMASS (Mahmood et al., 2019) skeleton representation with 22 joints, including the root joint. Each joint is represented by $\{p_{g},v_{g},r_{6d}\}$, where $p_{g}\in\mathbb{R}^{3}$ is the global position, $v_{g}\in\mathbb{R}^{3}$ is the global velocity, and $r_{6d}\in\mathbb{R}^{6}$ is the local 6D rotation of each joint, rendering $\mathbf{m}_{p}\in\mathbb{R}^{N\times 22\times 12}$. InterX follows the SMPL-X (Pavlakos et al., 2019) skeleton representation, comprising 54 body, hands and face joints, accompanied by root orientation and translation. Each joint and root orientation is represented by $\{r_{6d}\}$ and root translation by $\{p_{g}\}$, to which we incorporate root velocity rendering root $\{p_{g},v_{g}\}$ and $\mathbf{m}_{p}\in\mathbb{R}^{N\times 56\times 6}$. We adhere to the respective body representations of both datasets to demonstrate that our framework is compatible with different representation formats and joint counts. Other implementation details including architecture, training and inference hyper-parameters are provided in appendix E.
@@ 4.1 Comparison with state-of-the-art approaches
Quantitative Comparison Table 1 shows quantitative comparison of our InterMask with previous methods. All evaluations are run 20 times (with the exception of MModality, which is run 5 times), and we report the averaged results along with a 95% confidence interval. InterMask achieves state-of-the-art (SOTA) results on both InterHuman and InterX datasets. It records the lowest FID scores (5.154 on InterHuman and 0.399 on InterX), indicating superior realism and quality of generated interactions, and leads in R-Precision and MMdist, showing excellent text alignment. With different body representations in each dataset, these results highlight the robustness of our method, proving it is not dependent on a specific body pose structure. Furthermore, the larger performance gap on the larger InterX dataset suggests that InterMask scales effectively with more data. While our MModality is slightly lower than some methods, the high R-Precision and low MMdist emphasize that InterMask prioritizes adherence to text over extreme diversity, yet still achieves high Diversity scores, demonstrating its ability to generate a broad range of interactions. Please refer to appendix F for visual demonstration of InterMask’s ability to generate diverse results from the same text prompt.
Qualitative Comparison In Figure 4, we provide a qualitative comparison of interaction sequences generated by our InterMask and InterGen Liang et al. (2024), trained on the InterHuman dataset, for the same text descriptions. For the prompt ”The first person raises the right leg aggressively towards the second”, InterGen incorrectly raises both legs, while InterMask accurately raises only the right leg. For ”Two people bow to each other”, InterGen introduces an unnecessary fighting gesture after the bow, indicating overfitting to implicit biases in the training data, whereas InterMask stays true to the simple bow. Lastly, for ”One person is sitting and waving their hands at the other person, while the other drifts away”, InterGen generates a crouching pose instead of sitting and misplaces the waving action, while InterMask faithfully generates the described scene. These examples demonstrate that InterMask produces more higher quality and more reliable interactions compared to InterGen. For more qualitative results please refer to appendix A.
@@
User Study User studies are essential in generative AI, complementing metrics like FID by incorporating human perception to evaluate the realism and quality of generated content. To compare our InterMask with InterGen (Liang et al., 2024), we conducted a user study with 16 participants, where 30 interaction sequences were generated by both models with identical motion lengths. Each sequence pair was rated by 3 distinct users on interaction quality and adherence to text (scores out of 5), and their preferred interaction sequence among the two. More details are provided in appendix K. As shown in Figure 5, InterMask outperforms InterGen in interaction quality (4.089 vs. 3.296) and text adherence (3.938 vs. 3.198), with 69.14% of users preferring InterMask over InterGen.
@@ 4.2 Ablation Studies
Table 2 presents an ablation study to evaluate key components of our proposed Motion VQ-VAE. We investigated two primary factors: the input modality (individual motions vs two-person interactions) and the token map dimensions (1d vs 2d), using Reconstruction FID and MPJPE (Mean Per Joint Position Error). The results demonstrate that using individual motions as input with a 2D token map outperforms both the baseline of individual motions with a 1D token map, and the two-person interaction with a 2D token map. These findings suggest that preserving spatial information through 2D token maps and focusing on individual motions performs accurate motion reconstruction, and creates a more expressive token map with high-fidelity movement details.
Table 3 presents an ablation study to evaluate key components of our Inter-M Transformer. We compared collaborative modeling against alternative modeling and examined the impact of various attention mechanisms and masking strategies. Alternative Modeling (appendix G) models tokens of one person at a time, while conditioning in on the tokens of the other person, leading to an alternative generation during inference. Collaborative Modeling on the other hand, models tokens of both individuals simultaneously. Our full InterMask model with collaborataive modeling outperforms all ablated versions, particularly in R-Precision and FID scores. Notably, removing any of the attention mechanisms or employing alternative modeling severely degraded performance, with spatio-temporal attention being the most crucial one. Additionally, step unroll and interaction masking proved to be further fine-tuning the results. These results highlight the synergistic importance of all components in our transformer architecture for effective interaction modeling. Qualitative results for selected ablation studies are shown in appendix H.
@@ 4.3 Application: Reaction Generation
One of the key strengths of our InterMask model is its seamless support for the task of reaction generation with or without a text condition, where the motion of one person in an interaction is generated based on the provided reference motion of the other person. Unlike previous diffusion-based models, such as InterGen (Liang et al., 2024), InterMask can perform this task without any task-specific fine-tuning or architectural re-design. This flexibility arises from our use of the masked modeling technique combined with an individual motion tokenizer, which does not require interactions as input. During inference, we simply tokenize the reference motion, leave its tokens unmasked, and progressively generate tokens of the other person. More details are provided in appendix I.